# Recommendation and Temptation

### v1

@@ 4.2. Estimation Framework
    The set of items, $S_{j}(t)$, recommended to each user $j$ in each round $t$, along with the user’s subsequent choice. Notably, we assume the platform can identify when a user opts for an outside option over recommended items, but cannot determine which specific outside option was selected.
@@
    The explicit feedback or rating $r_{ij}$ provided by user $j$ after consuming item $i$. Crucially, we assume the platform can compute the feedback score from the provided rating, and vice versa.
@@
    We posit that the choice scores of outside options adhere to a known distributional form. This allows the platform to estimate the distribution parameters from available data, enabling the recommendation algorithm to compute outside option choice scores without requiring detailed knowledge of each user’s specific outside options. For our purposes, we assume these choice scores follow a normal distribution $\mathcal{N}(\mu,\sigma^{2})$, and we only need to estimate $\mu$ and $\sigma$.
@@
These assumptions, while providing crucial information, remain reasonable for adoption by real-world platforms. They offer the minimum necessary data for our proposed recommendation strategy to function effectively. Given these assumptions, we can solve for user-specific, item-specific and outside option-specific parameters. The detailed optimization process for this estimation is outlined in Algorithm 1.
@@
       for appropriate weights $\alpha$, $\beta$ with $\alpha+\beta=1$
      Compute $\hat{u}_{j}(i)=\hat{\mathbf{a}}^{\top}_{j}\hat{\mathbf{x}}_{i},\hat{v}_{j}(i)=\hat{\mathbf{b}}^{\top}_{j}\hat{\mathbf{y}}_{i}$;
       Return $\hat{u}_{j}(i),\hat{v}_{j}(i),\hat{\lambda}^{F}_{j},\hat{\lambda}^{C}_{j},\mu,\sigma$ for all $j,i$;
@@ 5. Evidence from Simulation
    Perfect information: The platform has complete knowledge of the enrichment and temptation values for all on-platform items and off-platform outside options for each user, the probability distribution governing the availability of outside options, and the user-specific feedback and choice parameters. The only unknown in this scenario is the specific outside option available in each round.
@@ 5.1. Results
To assess the efficacy of recommendation algorithms in ensuring user enrichment from consumption, we introduce the metric of overall individual enrichment. This measure quantifies the total enrichment an individual user receives, on average, from 50 rounds of consumption across both on-platform and off-platform options, following the initial warm-up period. This comprehensive metric allows us to evaluate how well each algorithm balances the immediate appeal of content with its long-term value to users.
@@
Figure 1(a) shows the overall individual enrichment when on-platform items offer higher enrichment and lower temptation compared to outside options. Traditional recommenders falter here by potentially suggesting low-temptation items that fail to prevent users from choosing low-enrichment outside options. For example, if a user’s outside option is online gambling, recommending an informational video on large language models is unlikely to be effective. Our algorithm addresses this challenge by recommending high-temptation on-platform items that still provide better enrichment than the outside option. In the gambling example, our algorithm might suggest a pop science video such as MythBusters, keeping the user engaged on the platform and away from gambling while also providing enrichment. Notably, click-based and purely temptation-based algorithms outperform ratings-based or purely enrichment-based ones in this scenario as well, simply by keeping users engaged on the platform. In the gambling example, these algorithms are likely recommending cat videos to users, which do not have much enrichment but keeps the users away from gambling regardless.
Conversely, figure 1(b) depicts the overall individual enrichment when on-platform items offer lower enrichment and higher temptation compared to outside options. In this scenario, traditional recommenders (e.g., ratings-based or click-based) primarily diminish user enrichment by suggesting on-platform items with higher temptation, thereby deterring users from selecting more enriching outside options. For instance, a video streaming platform might keep a student from exam preparation by recommending cat videos. Purely enrichment-based recommendations also fail to address this issue, as the outside option could surpass all on-platform items in enrichment while falling short in choice score. Our proposed algorithm, recognizing this complex dynamic, recommends items that guide users towards the optimal enrichment, potentially outside the platform. In the student example, our algorithm might suggest an informational video on large language models, likely prompting the student to return to exam preparation.
@@ 7. Recommendation Design Philosophy
(a) Frequency of consumption against enrichment of items
(b) Frequency of consumption against temptation of items
@@
Traditional recommender systems have predominantly adopted a predictive role in understanding and guiding user behavior on online platforms, often leaving the responsibility of self-control and consumer well-being entirely to users themselves. This approach, while seemingly neutral, can inadvertently lead to suboptimal outcomes for users, particularly when confronted with the reality of temptation in content consumption. For example, as we have remarked before in section 3.1, offering users more freedom of choice cannot hurt in a world without temptation, because users could simply not choose the additional recommended item(s). However, in the presence of temptation, the additional recommended item(s) could have high temptation and low enrichment, hindering users from choosing an item that would otherwise offer better enrichment. Traditional recommenders do not account for this complex dynamics, leaving everything up to the users and therefore providing little help in overcoming temptation. Our proposed recommendation algorithm diverges from this conventional stance, instead adopting a more prescriptive role that actively seeks to guide users towards consumption with higher enrichment. This shift in approach is grounded in the concept of libertarian paternalism, as conceptualized by Sunstein and Thaler (Sunstein and Thaler, 2003), which posits that private and public institutions should influence people’s behavior and guide them towards better versions of themselves while still respecting their freedom of choice. The existence of temptation and its profound effects on consumption choices makes this philosophy even more crucial for recommender systems. As demonstrated by the performance of baseline algorithms in our results, taking a purely predictive stance and failing to actively counter the effects of temptation when recommending items can lead to substantial harm to users. Our approach, therefore, represents a more responsible and user-centric model of recommendation, one that balances immediate user preferences with long-term user benefits.
While advocating for a more prescriptive approach, we simultaneously recognize the importance of stakeholder participation in defining and achieving social good. This belief is reflected in our recommendation design, which, in line with suggestions from previous research such as Ekstrand and Willemsen (Ekstrand and Willemsen, 2016), benefits from user participation beyond mere historical consumption data. Specifically, integrating user feedback about their outside options could improve our algorithm’s estimates of enrichment and temptation, thereby improving recommendation performance. This approach could not only enhance the algorithm’s effectiveness but also provide a channel for users to actively shape their digital experiences. We argue that platforms should not only consider but actively elicit such feedback from their users. This practice would enable platforms to gain a deeper understanding of their impact on user experience and provide a mechanism for continuous improvement and adaptation. By fostering this two-way communication, platforms can create a more transparent and user-centric environment, potentially leading to increased trust and satisfaction among users.
@@ 8. Conclusion and Future Work
In this paper, we addressed a key limitation of classic recommender systems: the assumption that users are enrichment maximizers, despite often succumbing to the temptation of instant gratification. We proposed a novel model of user consumption behavior that disentangles enrichment and temptation, and developed an optimal recommendation algorithm to maximize user enrichment under perfect information. Our approach identifies necessary information beyond historical consumption data and outlines a basic estimation framework using minimal assumptions. Through synthetic simulations and simulations based on real-world data from MovieLens, we demonstrated that our proposed algorithm has the potential to outperform traditional recommenders, providing users with superior enrichment from their consumption both on and off the platform. We note that if our estimation procedure is improved, this may lead to additional performance gains for our proposed recommendation strategy.
Our work opens several promising avenues for future research. First, improving the accuracy of our estimation framework could bring the algorithm’s performance closer to its theoretical optimum under perfect information. Second, further investigation into data collection methods and the validity of our assumptions would be invaluable for the practical implementation of our algorithm. Extending the model to consider multiple competing platforms and the dynamics of user behavior in a multi-platform environment represents another exciting direction. This could involve exploring how users navigate between platforms and how their preferences evolve in response to recommendations across different platforms.
Additionally, incorporating additional insights from behavioral economics, psychology, and related fields could lead to the development of more sophisticated and realistic models of user behavior. This interdisciplinary approach has the potential to create recommendation systems that not only predict user preferences more accurately but also actively contribute to user well-being.
@@
Appendix A Proof of Theorem 3.1
@@
Let $\mathcal{A}$ be the optimal recommendation strategy for $T+1$ rounds, and let $\mathcal{G}$ be the locally greedy one. Let $\mathcal{T}(\mathcal{A})$ and $\mathcal{T}(\mathcal{G})$ be the corresponding tree representations of these strategies. Let $A\rightarrow_{Y}B$ denote an edge in these trees where item $A$ was recommended and consumed, and then item $B$ was recommended. Define $A\rightarrow_{N}B$ analogously. Note that we can write any path in these trees as a collection of edges from item to item. Let $i_{k}$ denote the $k$-th new item the locally greedy strategy would recommend.
@@ Appendix B Simulation Parameter Selection
    The first components (representing universal enrichment/temptation) are drawn from a a normal distribution and are strongly anti-correlated: for all items $i=1,\ldots,n$,

### v2

@@ 4.2. Estimation Framework
    Behavioral data: The set of items, $S_{j}(t)$, recommended to each user $j$ in each round $t$, along with the user’s subsequent choice. Notably, we assume the platform can identify when a user opts for an outside option over recommended items, but cannot determine which specific outside option was selected.
@@
    Explicit feedback: Rating $r_{ij}$ provided by user $j$ after consuming item $i$. Crucially, we assume the platform can compute the feedback score from the provided rating, and vice versa.
@@
    We posit that the choice scores of outside options adhere to a known distributional form. This allows the platform to estimate the distribution parameters from available data, enabling the recommendation algorithm to compute outside option choice scores without requiring detailed knowledge of each user’s specific outside options. For our purposes, we assume these choice scores follow a normal distribution $\mathcal{N}(\mu,\sigma^{2})$, and we only need to estimate $\mu$ and $\sigma$. However, note that our framework could easily be extended to other distributions.
@@
These assumptions, while providing crucial information, remain reasonable for adoption by real-world platforms. They offer the minimum necessary data for our proposed recommendation strategy to function effectively. Given these assumptions, we can solve for user-specific, item-specific and outside option-specific parameters. The detailed optimization process for this estimation is outlined in Algorithm 1.  [^2] Code used for the experiments available at https://github.com/Sanzeed/recommendation-and-temptation
@@
6       for appropriate weights $\alpha$, $\beta$ with $\alpha+\beta=1$
7      Compute $\hat{u}_{j}(i)=\hat{\mathbf{a}}^{\top}_{j}\hat{\mathbf{x}}_{i},\hat{v}_{j}(i)=\hat{\mathbf{b}}^{\top}_{j}\hat{\mathbf{y}}_{i}$;
8       Return $\hat{u}_{j}(i),\hat{v}_{j}(i),\hat{\lambda}^{F}_{j},\hat{\lambda}^{C}_{j},\mu,\sigma$ for all $j,i$;
9
10
@@ 5. Evidence from Simulation
    Perfect information: The platform has complete knowledge of the enrichment and temptation of all on-platform items and off-platform outside options for each user, the probability distribution governing the availability of outside options, and the user-specific feedback and choice parameters. The only unknown in this scenario is the specific outside option available in each round.
@@ 5.1. Results
To assess the efficacy of recommendation algorithms in ensuring user enrichment from consumption, we introduce the metric of overall individual enrichment. This measure quantifies the total enrichment an individual user receives, on average, from 50 rounds of consumption across both on-platform and off-platform options, following the initial warm-up period. This comprehensive metric allows us to evaluate how well each algorithm balances the immediate appeal of content with its inherent value to a user.
@@
Figure 1(a) shows the overall individual enrichment when on-platform items offer higher enrichment and lower temptation compared to outside options. Traditional recommenders falter here by potentially suggesting low-temptation items that fail to prevent users from choosing low-enrichment outside options. Consider the example where a user’s outside option is online gambling. Recommending an informational video on large language models in this case is unlikely to be effective. Our algorithm addresses this challenge by recommending high-temptation on-platform items that still provide better enrichment than the outside option. In the gambling example, our algorithm might suggest a pop science video such as MythBusters, keeping the user engaged on the platform and away from gambling while also providing enrichment. Notably, click-based and purely temptation-based algorithms outperform ratings-based or purely enrichment-based ones in this scenario as well, simply by keeping users engaged on the platform. In the gambling example, these algorithms are likely recommending cat videos to users, which do not have much enrichment but keeps the users away from gambling regardless.
Conversely, figure 1(b) depicts the overall individual enrichment when on-platform items offer lower enrichment and higher temptation compared to outside options. In this scenario, traditional recommenders (e.g., ratings-based or click-based) primarily diminish user enrichment by suggesting on-platform items with higher temptation, thereby deterring users from selecting more enriching outside options. Consider the example where a video streaming platform might keep a student from exam preparation by recommending cat videos. Purely enrichment-based recommendations also fail to address this issue, as the outside option could surpass all on-platform items in enrichment while falling short in choice score. Our proposed algorithm, recognizing this complex dynamic, recommends items that guide users towards the optimal enrichment, potentially outside the platform. In the student example, our algorithm might suggest an informational video on large language models, likely prompting the student to return to exam preparation.
@@ 7. Recommendation Design Philosophy
Traditional recommender systems have predominantly adopted a predictive role in understanding and guiding user behavior on online platforms, often leaving the responsibility of self-control and consumer well-being entirely to users themselves. This approach, while seemingly neutral, can inadvertently lead to suboptimal outcomes for users, particularly when confronted with the reality of temptation in content consumption. For example, as we have remarked before in section 3, offering users more “freedom” of choice cannot hurt in a world without temptation, because users could simply not choose the additional recommended item(s). However, in the presence of temptation, the additional recommended item(s) could have high temptation and low enrichment, hindering users from choosing an item that would otherwise offer better enrichment.
The act of recommending tempting but unenriching content can be conceptualized as creating a form of harm: it generates desires that are costly to fulfill yet unsatisfying to leave unmet. This reveals a subtle challenge for recommender systems, as eliciting engagement signals from users may be more costly than previously recognized. Ideally, platforms seek to learn user preferences by recommending diverse sets of items (Peng et al., 2023). However, larger option pools are more likely to contain highly tempting options that deter users from selecting more enriching alternatives. This creates a fundamental trade-off for recommenders between maximizing user enrichment and gathering information about user preferences.
Traditional recommenders do not account for the complex dynamics of temptation and enrichment, leaving everything up to the users and therefore providing little help in overcoming temptation. Our proposed recommendation algorithm diverges from this predictive stance, instead adopting a more prescriptive role that actively seeks to guide users towards consumption with higher enrichment. This shift in approach is similar in spirit to the concept of libertarian paternalism, as conceptualized by Sunstein and Thaler (Sunstein and Thaler, 2003), which posits that private and public institutions should guide people towards better versions of themselves while still respecting their freedom of choice. Crucially, note that the role of our proposed recommender is not a “pro-social” one —we are not imposing any moral constraints beyond users’ underlying preferences or desires via our recommender. Instead, we merely uncover user preferences or desires otherwise obscured by temptation, and provide better navigation towards them. Any system must wrestle with the choice of whether to optimize for user’s perceived enrichment or the user’s perceived desires under the revealed preference assumption. This is not really a “paternalistic" a choice of any kind, rather, in the conception of the dual self, an choice of which “self” to serve. Perhaps users could, at the beginning, be offered a choice between our recommendation system and an engagement-based approach. In the dual self conception of temptation, the users would always choose our system because it will better help to achieve the goals the “self" that is doing the choosing.
@@ 8. Conclusion and Future Work
In this paper, we addressed a key limitation of classic recommender systems: the assumption that users are enrichment maximizers, despite often succumbing to the temptation of instant gratification. We introduced a novel model of user consumption behavior that disentangles enrichment and temptation and accounts for the existence of outside options. We highlight the challenge in recognizing what to optimize for when recommending content, proposed the appropriate objective for enrichment maximization, and developed the optimal strategy under perfect information. Our approach identifies the information about user experience outside the platform necessary beyond historical consumption data on the platform, and outlines a basic estimation framework using minimal assumptions. Through synthetic simulations and simulations based on real-world data from MovieLens, we demonstrated that our proposed algorithm has the potential to outperform traditional recommenders, providing users with superior enrichment from their consumption both on and off the platform. We note that if our estimation procedure is improved, this may lead to additional performance gains for our proposed recommendation strategy.
Our work opens several promising avenues for future research. First, improving the accuracy of our estimation framework could bring the algorithm’s performance closer to its theoretical optimum under perfect information. Second, further investigation into data collection methods, specifically eliciting information about outside options, and the validity of our assumptions would be invaluable for the practical implementation of our algorithm. Extending the model to consider multiple competing platforms and the dynamics of user behavior in a multi-platform environment represents another exciting direction. This could involve exploring how users navigate between platforms and how their preferences evolve in response to recommendations across different platforms.
@@
Appendix A Proof of Theorem 1
Theorem 1.
@@ Appendix A Proof of Theorem 1
Let $\mathcal{A}$ be the optimal recommendation strategy for $T+1$ rounds, and let $\mathcal{G}$ be the locally greedy one. Let $\mathcal{T}(\mathcal{A})$ and $\mathcal{T}(\mathcal{G})$ be the corresponding tree representations of these strategies. Let $A\rightarrow_{Y}B$ denote an edge in these trees where item $A$ was recommended and consumed, and then item $B$ was recommended. Define $A\rightarrow_{N}B$ analogously. Note that we can write any path in these trees as a collection of edges from item to item. Let $i_{k}$ denote the $k$th new item the locally greedy strategy would recommend.
@@ Appendix B Simulation Parameter Selection
Figure 3. Overall individual enrichment provided by recommendation algorithms when enrichment and temptation of on-platform items have skewed, fat-tailed distributions. (a) shows algorithm performance for enriching on-platform items, while (b) shows performance for tempting on-platform items. The striped bars show performance when the platform has partial information only. Our proposed algorithm maintains superior performance.
@@
    The first components (representing universal enrichment and temptation) are drawn from a a normal distribution and are strongly anti-correlated: for all items $i=1,\ldots,n$,
@@
Appendix C Robustness Tests
In realistic scenarios, enrichment and temptation of items for a given user may not be normally distributed. To ensure that our estimation framework is robust to the distributions of these quantities, we reran our simulated experiments by drawing the first components of representations $x_{i1},y_{i1}$ of each on-platform item $i$ (representing universal enrichment and temptation) from skewed, fat-tailed Johnson’s $S_{u}$ distributions. Specifically,
$x_{i1}$ $\sim Johnson-S_{U}\left(3.25,1,12.3520,0.3933\right)$
$y_{i1}$ $\sim Johnson-S_{U}\left(3.25,1,2.3520,0.3933\right)$
We kept the distribution of enrichment and temptation of outside options identical to our previous experiments.
Figure 3 shows the performance of our algorithm against four baseline algorithms in four possible scenarios described in section 5. As the figure shows, our algorithm maintains its superior performance, demonstrating its robustness to distributions of enrichment and temptation.